OpenStudy (brynndelyn):

Which of the following is a solution of x^2+5x=-2? 5 plus or minus the square root of 33 divided by two. 5 plus or minus the square root of 17 divided by two negative 5 plus or minus the square root of 33 divided by two negative 5 plus or minus the square root of 17 divided by two

OpenStudy (anonymous):

you have to use the quadratic equation. do you know how?

OpenStudy (brynndelyn):

Sorta

OpenStudy (campbell_st):

well rewrite the equation as x^2 + 5x + 2 = 0 you need the general quadratic formula a = 1, b = 5 and c = 1 the formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] hope it helps

OpenStudy (campbell_st):

oops c = 2...

OpenStudy (isaiah.feynman):

Or you can complete the square.